OpenStudy (anonymous):

"The straight line equation y=(3/4)x makes an acute angle k with the x-axis. Find the value of sin2k and cos2k"

OpenStudy (sanra123):

can u help me plz

OpenStudy (anonymous):

I'm sorry, I can't!

OpenStudy (sanra123):

why?

OpenStudy (anonymous):

Because I don't know how to work it out.

OpenStudy (sanra123):

ok thx

OpenStudy (dumbcow):

in general, for any line the slope equals the tangent \[\tan (k) = slope = \frac{3}{4}\] find sin(k), cos(k) using right triangle  notice i solved for hypotenuse using pythagorean thm now using SOH CAH TOA \[\sin(k) = \frac{3}{5}\] \[\cos(k) = \frac{4}{5}\] double angle identities \[\sin (2k) = 2\sin k \cos k = 2*\frac{3}{5}*\frac{4}{5} = \frac{24}{25}\] \[\cos (2k) = \cos^{2} k - \sin^{2} k = (\frac{4}{5})^{2} - (\frac{3}{5})^{2} = \frac{7}{25}\] there you go, hope that helps